# Unifying Bias and Unfairness in Information Retrieval: A Survey of Challenges and Opportunities with Large Language Models

### v1

@@ Abstract.
With the rapid advancement of large language models (LLMs), information retrieval (IR) systems, such as search engines and recommender systems, have undergone a significant paradigm shift. This evolution, while heralding new opportunities, introduces emerging challenges, particularly in terms of biases and unfairness, which may threaten the information ecosystem. In this paper, we present a comprehensive survey of existing works on emerging and pressing bias and unfairness issues in IR systems when the integration of LLMs. We first unify bias and unfairness issues as distribution mismatch problems, providing a groundwork for categorizing various mitigation strategies through distribution alignment. Subsequently, we systematically delve into the specific bias and unfairness issues arising from three critical stages of LLMs integration into IR systems: data collection, model development, and result evaluation. In doing so, we meticulously review and analyze recent literature, focusing on the definitions, characteristics, and corresponding mitigation strategies associated with these issues. Finally, we identify and highlight some open problems and challenges for future work, aiming to inspire researchers and stakeholders in the IR field and beyond to better understand and mitigate bias and unfairness issues of IR in this LLM era. We also consistently maintain a GitHub repository for the relevant papers and resources in this rising direction at https://github.com/KID-22/LLM-IR-Bias-Fairness-Survey.
Bias, Fairness, Information Retrieval, Large Language Models
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/xxxxxxxconference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Information retrievalccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. Overview of three stages of the intersection between LLMs and IR systems. (a) LLMs-generated content as new IR data sources. (b) Incorporating LLMs to enhance or as IR models. (c) Adopting LLMs as results evaluators in IR systems.
Information Retrieval (IR) systems strive to navigate the era of information overload, facilitating users in acquiring information more efficiently and effectively (Xu et al., 2018; Singhal et al., 2001; Manning, 2009). The integration of Large Language Models (LLMs) has fundamentally redefined IR systems, including the introduction of LLM-generated data as new IR data sources, the shift from passive retrieval to proactive generation as core paradigms, and adopting LLMs as results evaluators for IR systems (Zhu et al., 2023; Ai et al., 2023). These advancements, however, bring forth new challenges in bias and unfairness, affecting the reliability of IR systems and potentially contributing to societal issues like echo chambers (Del Vicario et al., 2016; Sharma et al., 2024) and cognitive interference (Sarason et al., 2014; Mihaylov et al., 2018). For instance, researchers found that LLMs often retrieve information that deviates from facts and is biased towards LLM-generated content (Dai et al., 2023b; Xu et al., 2023b; Li et al., 2023b; Chern et al., 2023). Moreover, LLMs frequently manifest stereotypes and discriminatory content to users and amplify disparities between items of different socio-economic statuses (Zhang et al., 2023a; Xu et al., 2023e; Jiang et al., 2024a).
@@
Our survey begins with a brief overview of how LLMs are integrated into IR systems, setting the stage for understanding the emergence of new bias and unfairness challenges across three pivotal stages of the IR lifecycle: data collection, model development, and result evaluation. Then we propose a unified perspective on bias and unfairness, categorizing them as distribution mismatch problems. Based on this unified view, we categorize mitigation strategies into two principal groups: data sampling, including data augmentation and data filtering, and distribution reconstruction, encompassing rebalancing, regularization, and prompting. Following this taxonomy, we delve into a detailed analysis of several distinct types of bias and unfairness phenomena that arise with the integration of LLMs into IR systems, spanning the aforementioned stages. Our systematic review encompasses a comprehensive examination of these issues and their respective mitigation strategies in recent studies, providing a holistic view of the current landscape and guiding future efforts in eliminating bias and ensuring fairness for more trustworthy IR systems.
@@
In summary, the key contributions of this work are as follows:
$\bullet$ We provide a novel unified perspective for understanding bias and unfairness as distribution mismatch problems, alongside a detailed review of several types of bias and unfairness arising from the integration of LLMs into IR systems.
$\bullet$ We systematically organize mitigation strategies into two key categories: data sampling and distribution reconstruction, offering a comprehensive roadmap for effectively combating bias and unfairness with state-of-the-art approaches.
$\bullet$ We identify the current challenges and future directions, providing insights to facilitate the development of this potential and demanding research area.
Organization of Our Survey. The remaining part of this survey is organized as follows: Section 2 introduces the utilization of the distribution mismatch perspective to unify bias and unfairness in IR with LLMs. Section 3 and Section 4 illustrate the cause and the mitigation strategies for the new or significantly evolved issues related to bias and unfairness in IR with LLMs. Section 5 outlines the future challenges and opportunities associated with bias and unfairness issues in IR with LLMs.
@@ 2. A Unified View of Bias and Unfairness
In this section, we first introduce the background of integrating LLMs into IR systems, where new issues related to bias and unfairness emerge. Then we provide a unified view of bias and unfairness from the perspective of distribution alignment and outline the mitigation strategies based on this view.
@@ 2.1. Background
LLMs-generated content as new IR data sources. The emergence of LLMs has significantly accelerated the growth of Artificial Intelligence Generated Content (AIGC), marking a new era in content creation. Unlike traditional Professional and User Generated Content (PGC and UGC) sources, AIGC can be produced automatically at scale, potentially dominating the content landscape. (Cao et al., 2023; Wu et al., 2023a; Hanley and Durumeric, 2023). However, AIGC also reshapes the distribution of the IR data, resulting in new concerns about bias and fairness.
Incorporating LLMs to enhance or as IR models. The impressive emergent capabilities of LLMs in understanding, reasoning, and generalization have motivated significant efforts to integrate them into the development of next-generation IR systems (Zhu et al., 2023). On one hand, LLMs have been deployed to refine key components of traditional IR systems (Sun et al., 2023; Shi et al., 2023b; Wang et al., 2023g), enhancing their effectiveness and efficiency. On the one hand, beyond enhancing existing frameworks, LLMs also introduce a novel paradigm by acting as generative search and recommendation agents (Nakano et al., 2021; Gur et al., 2023; Dai et al., 2023a; Zhang et al., 2023e), directly generating responses to fulfill user queries.
Adopting LLMs as results evaluators in IR systems. Human evaluation plays a pivotal role in IR systems, particularly in conversational search and recommendation, such as directly assessing the quality of responses from generative LLM-based IR models (Dai et al., 2023a; Li et al., 2023h; Zheng et al., 2023b). However, human evaluation comes with significant challenges, including high costs and a lack of reproducibility. LLMs, with their advanced language modeling and understanding capabilities, offer new possibilities for conducting evaluations of these complex tasks without human evaluation (Chiang and Lee, 2023; Svikhnushina and Pu, 2023; Li et al., 2024; Gao et al., 2024). This shift not only streamlines the evaluation process but also mitigates the substantial costs associated with human evaluations, further facilitating the development of IR systems.
While the reshaping of the above IR stages by LLMs has introduced numerous new opportunities, it has also given rise to many new and pressing issues related to bias and unfairness in these stages during the integration.
@@ 2.2. Distribution Alignment Perspective
In this section, we formulate the problems of bias and fairness from a distribution alignment perspective, offering a unified framework for understanding these challenges and inspiring mitigation strategies.
Formally, when a user interacts with a typical IR system, she may optionally provide her personalized information requirement (commonly represented as text-based requirements $T$) along with their personalized attributes (typically indicated as user profile $U$) to the system. Subsequently, the goal of the IR system is to retrieve the target information $R$ (e.g., documents, items, or advertisements, et al. ) for this user, with the user’s information requirements and optional interaction history $H$ as the input $Q=\{T,U,H\}$. Let $\widehat{R}=f(Q)$ be the predicted result either from existing IR models or directly generated by LLMs, where $f(\cdot)$ is the model.
Let $P(\widehat{R})$ be the distribution of predicted results for all users, we can unify the bias and unfairness problem as a distribution mismatch problem, which be can formally defined as follows:
(1) $P(\widehat{R})\neq P(R),$
where $P(R)$ is the ground truth distribution of the target results.
@@ 3. Cause and Mitigation of Bias
Data Collection Source Bias (Dai et al., 2023b; Xu et al., 2023b)
Factuality Bias (Ram et al., 2023; Yu et al., 2023b; Xu et al., 2023c; Press et al., 2022; Xu et al., 2024b, a; Gunasekar et al., 2023) (Gunasekar et al., 2023; Touvron et al., 2023; Yang et al., 2022) (Xu et al., 2023c; Press et al., 2022; Wang et al., 2023f; Sun et al., 2022a)
@@
Result Evaluation Selection Bias (Li et al., 2023e; Zheng et al., 2023b; Pezeshkpour and Hruschka, 2023; Chern et al., 2024; Zheng et al., 2024; Li et al., 2023d; Wang et al., 2023b; Chu et al., 2024; Yinhong et al., 2024) (Zheng et al., 2023c; Wang et al., 2023b; Liusie et al., 2024) (Zheng et al., 2024; Pezeshkpour and Hruschka, 2023; Wang et al., 2023b; Kim et al., 2023)
@@ 3.1.1. Source Bias
Source Bias: IR models tend to rank content generated by LLMs higher than content authored by humans.
Specifically, as LLMs fuel the rapid expansion of LLM-generated content, the corpus for IR systems now increasingly encompasses a mix of human-written and LLM-generated texts. Recent studies (Dai et al., 2023b; Xu et al., 2023b) highlight that modern retrieval models, especially those leveraging neural matching techniques, tend to favor LLM-generated content over human-authored content with similar semantics. This preference stems from the unique representations embedded in LLM-generated content, which neural retrieval models can capture and thus assign a higher estimated relevancy score (Dai et al., 2023b; Xu et al., 2023b). More severely, this source bias is further amplified when LLM-generated content is included in model training, presenting a challenge to the information ecosystem. (Dai et al., 2023b; Xu et al., 2023b). Tan et al. (2024) further reveal that this bias will extend from retrievers to the readers.
To counteract source bias, recent studies have introduced debiased constraints into the training objectives of neural retrieval models (Dai et al., 2023b; Xu et al., 2023b). This strategy aims to correct the skewed relevancy predictions favoring LLM-generated content, ensuring fair treatment between human-written and LLM-generated content. By adopting a distribution alignment perspective, such mitigation efforts strive to recalibrate the neural models’ relevancy distribution towards an ideal state, where documents are judged equally based on their semantic content rather than their source. Despite these advancements, the existing strategies require further refinement for enhanced adaptability, highlighting a need for more robust mitigation methods in future research.
@@ 3.1.2. Factuality Bias
Factuality Bias: LLMs may produce content that does not align with recognized factual information of the real world.
Many studies have shown that LLMs are at risk of generating factual errors. For instance, TruthfulQA (Lin et al., 2021) has highlighted that language models generate many false answers that mimic popular misconceptions in question-answering tasks and have the potential to deceive humans. Besides, merely scaling up models is not promising for improving truthfulness, which means that LLMs still face challenges in generating factually correct content. Lee et al. (2022) show LLMs are susceptible to generating text with nonfactual in an open-ended generation because of the “uniform randomness” at every sampling step. FActScore (Min et al., 2023) finds that LLMs lag significantly behind humans in ensuring the factual consistency of long-form text generation. KoLA (Yu et al., 2023a) shows alignment harms LLMs’ memorization of world knowledge. Other studies (McKenna et al., 2023) reveal that LLMs also exhibit factual hallucination in natural language inference tasks. In addition to the above studies, many large-scale benchmarks (Li et al., 2023b; Chern et al., 2023; Chen et al., 2023b) also indicate that LLMs exhibit factuality bias in multi-task and multi-domain scenarios.
@@
To mitigate factuality bias, in the training of LLMs, some methods focus on providing high-quality and factually correct training data for LLMs (Gunasekar et al., 2023; Touvron et al., 2023). In the inference, previous studies can be divided into two categories. One is with the help of an external factual knowledge base such as retrieval-augmented generation (Ram et al., 2023; Yu et al., 2023b; Xu et al., 2023c; Press et al., 2022; Xu et al., 2024b, a; Gao et al., 2022). The other is to improve the ability of LLMs themselves such as Self-Consistency (Wang et al., 2023f) and Dola (Chuang et al., 2023).
@@ 3.2.1. Position Bias
Position Bias: LLM-based IR models tend to give preference to documents or items from specific input positions.
Recent works have highlighted that the order of candidate documents or items can significantly impact the performance of LLM-based IR models, while conventional IR models are often not affected by the changing of input orders (Xu et al., 2024c; Hou et al., 2024; Tang et al., 2023). For instance, LLM-based models often show a preference for content positioned at the beginning or end of a list, neglecting the contributions of items in the middle. This “lost in the middle” suggests that these LLM-based models may not fully utilize the context provided by items that don’t occupy prominent positions in the input sequence (Liu et al., 2024).
There are a number of works on mitigating position bias, and we can categorize them into three lines based on our distribution alignment framework. (1) Prompting: This approach involves carefully designed prompts to encourage the models to disregard the input’s order (Hou et al., 2024; Dai et al., 2023a). Nonetheless, due to LLMs’ prompt sensitivity, this requires precise and task-specific prompt engineering across various tasks and domains. (2) Data Augmentation: This approach has been explored in numerous studies, which is a form of data augmentation that involves random shuffling of candidates followed by aggregation to determine the final ranking. (Hou et al., 2024; Qin et al., 2023; Tang et al., 2023; Wu et al., 2023b; Luo et al., 2023; Zhang et al., 2023b). For instance, Tang et al. (2023) introduced a permutation self-consistency method, offering theoretical guarantees under certain conditions, enabling models to produce and aggregate potential outcomes from various candidate permutations, enhancing result stability. (3) Rebalancing: This method counteracts position bias by adjusting the prior distribution sensitive to positions. It recalibrates the model’s output, addressing the inherent bias towards item positions (Ma et al., 2023; Wu et al., 2023b).

### v2

# Bias and Unfairness in Information Retrieval Systems: New Challenges in the LLM Era

@@ Abstract.
With the rapid advancements of large language models (LLMs), information retrieval (IR) systems, such as search engines and recommender systems, have undergone a significant paradigm shift. This evolution, while heralding new opportunities, introduces emerging challenges, particularly in terms of biases and unfairness, which may threaten the information ecosystem. In this paper, we present a comprehensive survey of existing works on emerging and pressing bias and unfairness issues in IR systems when the integration of LLMs. We first unify bias and unfairness issues as distribution mismatch problems, providing a groundwork for categorizing various mitigation strategies through distribution alignment. Subsequently, we systematically delve into the specific bias and unfairness issues arising from three critical stages of LLMs integration into IR systems: data collection, model development, and result evaluation. In doing so, we meticulously review and analyze recent literature, focusing on the definitions, characteristics, and corresponding mitigation strategies associated with these issues. Finally, we identify and highlight some open problems and challenges for future work, aiming to inspire researchers and stakeholders in the IR field and beyond to better understand and mitigate bias and unfairness issues of IR in this LLM era. We also consistently maintain a GitHub repository for the relevant papers and resources in this rising direction at https://github.com/KID-22/LLM-IR-Bias-Fairness-Survey.
Information Retrieval, Large Language Model, Bias, Fairness
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671458isbn: 979-8-4007-0490-1/24/08ccs: Information systems Information retrieval
@@ 1. Introduction
Figure 1. Overview of three stages of the intersection between LLMs and IR systems. (a) LLMs-generated content as new data sources for IR. (b) Incorporating LLMs to enhance or as IR models. (c) Adopting LLMs as results evaluators in IR systems.
Information Retrieval (IR) systems strive to navigate the era of information overload, facilitating users in acquiring information more efficiently and effectively (Xu et al., 2018; Singhal et al., 2001; Manning, 2009). The integration of Large Language Models (LLMs) has fundamentally redefined IR systems, including the introduction of LLM-generated data as new IR data sources, the shift from passive retrieval to proactive generation as core paradigms, and adopting LLMs as results evaluators for IR systems (Zhu et al., 2023; Ai et al., 2023). These advancements, however, bring forth new challenges in bias and unfairness, affecting the reliability of IR systems and potentially contributing to societal issues like echo chambers (Del Vicario et al., 2016; Sharma et al., 2024) and cognitive interference (Sarason et al., 2014; Mihaylov et al., 2018). For instance, researchers found that LLMs often retrieve information that deviates from facts and is biased towards LLM-generated content (Dai et al., 2024b; Xu et al., 2023b; Li et al., 2023b; Chern et al., 2023). Moreover, LLMs frequently manifest stereotypes and discriminatory content to users and amplify disparities between items of different socio-economic statuses (Zhang et al., 2023a; Xu et al., 2023d; Jiang et al., 2024a).
@@
Our survey begins with a brief overview of how LLMs are integrated into IR systems, setting the stage for understanding the emergence of new bias and unfairness challenges across three pivotal stages of the IR lifecycle: data collection, model development, and result evaluation. Then we propose a unified perspective on bias and unfairness, categorizing them as distribution mismatch problems. Based on this unified view, we categorize mitigation strategies into two principal groups: data sampling, including data augmentation and data filtering, and distribution reconstruction, encompassing rebalancing, regularization, and prompting. Following this taxonomy, we delve into a detailed analysis of several types of bias and unfairness phenomena that arise with the integration of LLMs into IR systems, spanning the aforementioned stages. Our systematic review encompasses a comprehensive examination of these issues and their respective mitigation strategies in recent studies, providing a holistic view of the current landscape and guiding future efforts in eliminating bias and ensuring fairness for more trustworthy IR systems.
@@
Summary of Contributions. (1) We provide a novel unified perspective for understanding bias and unfairness as distribution mismatch problems, alongside a detailed review of several types of bias and unfairness arising from integrating LLMs into IR systems. (2) We systematically organize mitigation strategies into two key categories: data sampling and distribution reconstruction, offering a comprehensive roadmap for effectively combating bias and unfairness with state-of-the-art approaches. (3) We identify the current challenges and future directions, providing insights to facilitate the development of this potential and demanding research area.
@@ 2. A Unified View of Bias and Unfairness
In this section, we provide a unified view of bias and unfairness from the perspective of distribution alignment and outline the mitigation strategies based on this view.
@@ 2.1. Background
LLMs-generated content as new data sources for IR. The emergence of LLMs has significantly accelerated the growth of Artificial Intelligence Generated Content (AIGC), marking a new era in content creation. Unlike traditional Professional and User Generated Content (PGC and UGC) sources, AIGC can be produced automatically at scale, potentially dominating the content landscape. (Cao et al., 2023; Wu et al., 2023a; Hanley and Durumeric, 2023). However, AIGC also reshapes the distribution of the IR data, resulting in new concerns about bias and fairness.
Incorporating LLMs to enhance or as IR models. The impressive emergent capabilities of LLMs in understanding, reasoning, and generalization have motivated significant efforts to integrate them into the development of next-generation IR systems (Zhu et al., 2023). On one hand, LLMs have been deployed to refine key components of traditional IR systems (Sun et al., 2023; Shi and Min, 2023; Wang et al., 2023g), enhancing their effectiveness and efficiency. On the one hand, beyond enhancing existing frameworks, LLMs also introduce a novel paradigm by acting as generative search and recommendation agents (Nakano et al., 2021; Gur et al., 2023; Dai et al., 2023; Zhang et al., 2023e), directly generating responses to fulfill user queries.
Adopting LLMs as results evaluators in IR systems. Human evaluation plays a pivotal role in IR systems, particularly in conversational search and recommendation, such as directly assessing the quality of responses from generative LLM-based IR models (Dai et al., 2023; Li et al., 2023h; Zheng et al., 2023b). However, human evaluation comes with significant challenges, including high costs and a lack of reproducibility. LLMs, with their advanced language modeling and understanding capabilities, offer new possibilities for conducting evaluations of these complex tasks without human evaluation (Chiang and Lee, 2023; Svikhnushina and Pu, 2023; Li et al., 2024; Gao et al., 2024). This shift not only streamlines the evaluation process but also mitigates the substantial costs associated with human evaluations, further facilitating the development of IR systems.
@@ 2.2. Distribution Alignment Perspective
While the reshaping of the above IR stages by LLMs has introduced numerous new opportunities, it has also given rise to many new and pressing issues related to bias and unfairness. In this section, we formulate the problems of bias and fairness from a distribution alignment perspective, offering a unified framework for understanding these challenges and inspiring mitigation strategies.
Formally, when a user interacts with a typical IR system, he/she may optionally provide his/her personalized information requirement $T$ along with their personalized attributes (typically indicated as user profile $U$) to the system. Subsequently, the goal of the IR system is to retrieve the target information $R$ (e.g., documents, items, or advertisements, et al. ) for this user, with the user’s information requirements and optional interaction history $H$ as the input $Q=\{T,U,H\}$. Let $\widehat{R}=f(Q)$ be the predicted result either from IR models or directly generated by LLMs, where $f(\cdot)$ is the model.
Let $P(\widehat{R})$ be the distribution of predicted results for all users, we can unify the bias and unfairness problem as a distribution mismatch problem with the ground truth distribution $P(R)$, where $R$ is the target result, which can be defined as follows:
(1) $P(\widehat{R})\neq P(R).$
@@ 3. Cause and Mitigation of Bias
Data Collection Source Bias (Chen et al., 2024b) (Dai et al., 2024b; Xu et al., 2023b; Zhou et al., 2024)
Factuality Bias (Ram et al., 2023; Yu et al., 2023; Xu et al., 2024b; Press et al., 2022; Xu et al., 2024c, a; Gunasekar et al., 2023) (Gunasekar et al., 2023; Touvron et al., 2023; Yang et al., 2022) (Xu et al., 2024b; Press et al., 2022; Wang et al., 2023f; Sun et al., 2022a)
@@
Result Evaluation Selection Bias (Li et al., 2023e; Zheng et al., 2023b; Pezeshkpour and Hruschka, 2023; Chern et al., 2024; Zheng et al., 2024; Li et al., 2023d; Wang et al., 2023b; Chu et al., 2024) (Zheng et al., 2023c; Wang et al., 2023b; Liusie et al., 2024) (Zheng et al., 2024; Pezeshkpour and Hruschka, 2023; Wang et al., 2023b; Kim et al., 2023)
@@ 3.1.1. Source Bias
$\bullet$ Definition. Information retrieval models tend to rank content generated by LLMs higher than content authored by humans.
Specifically, as LLMs fuel the rapid expansion of LLM-generated content, the corpus for IR systems now increasingly encompasses a mix of human-written and LLM-generated texts. Recent studies (Dai et al., 2024b; Xu et al., 2023b; Dai et al., 2024a) highlight that modern retrieval models, especially those leveraging neural matching techniques, tend to favor LLM-generated content over human-authored content with similar semantics. This preference stems from the unique representations embedded in LLM-generated content, which neural retrieval models can capture and thus assign a higher estimated relevancy score (Dai et al., 2024b; Xu et al., 2023b). More severely, this source bias is further amplified when LLM-generated content is included in model training, presenting a challenge to the information ecosystem (Xu et al., 2023b). Zhou et al. (2024) further explores this escalation of source bias in user, data, and recommender systems feedback loop. Tan et al. (2024) further that this bias will extend from retrievers to the readers and Chen et al. (2024b) uncover this bias in the RAG systems.
To counteract source bias, recent studies have introduced debiased constraints into the training objectives of IR retrieval models (Dai et al., 2024b; Xu et al., 2023b; Zhou et al., 2024). This strategy aims to correct the skewed relevancy predictions favoring LLM-generated content, ensuring fair treatment between human-written and LLM-generated content. By adopting a distribution alignment perspective, such mitigation efforts strive to recalibrate the IR models’ relevancy distribution towards an ideal state, where documents are judged equally based on their semantic content rather than their source.
@@ 3.1.2. Factuality Bias
$\bullet$ Definition. LLMs may produce content that does not align with recognized factual information of the real world.
Many studies have shown that LLMs are at risk of generating factual errors. For instance, TruthfulQA (Lin et al., 2021) has highlighted that language models generate many false answers that mimic popular misconceptions in question-answering tasks and have the potential to deceive humans. Besides, merely scaling up models is not promising for improving truthfulness, which means that LLMs still face challenges in generating factually correct content. Lee et al. (2022) show LLMs are susceptible to generating text with nonfactual in an open-ended generation because of the “uniform randomness” at every sampling step. FActScore (Min et al., 2023) finds that LLMs lag significantly behind humans in ensuring the factual consistency of long-form text generation. Other studies (McKenna et al., 2023) reveal that LLMs also exhibit factual hallucination in natural language inference tasks. In addition to the above studies, many large-scale benchmarks (Li et al., 2023b; Chern et al., 2023; Chen et al., 2023b) also indicate that LLMs exhibit factuality bias in multi-task and multi-domain scenarios.
@@
To mitigate factuality bias, in the training of LLMs, some methods focus on providing high-quality and factually correct training data for LLMs (Gunasekar et al., 2023; Touvron et al., 2023). In the inference, previous studies can be divided into two categories. One is with the help of an external factual knowledge base such as retrieval-augmented generation (Ram et al., 2023; Yu et al., 2023; Xu et al., 2024b; Press et al., 2022; Xu et al., 2024c, a; Gao et al., 2022). The other is to improve the ability of LLMs themselves such as Self-Consistency (Wang et al., 2023f) and Dola (Chuang et al., 2023).
@@ 3.2.1. Position Bias
$\bullet$ Definition. LLM-based IR models tend to give preference to documents or items from specific input positions.
Recent works have highlighted that the order of candidate documents or items can significantly impact the performance of LLM-based IR models, while conventional IR models are often not affected by the changing of input orders (Xu et al., 2024d; Hou et al., 2024; Tang et al., 2023). For instance, LLM-based models often show a preference for content positioned at the beginning or end of a list, neglecting the contributions of items in the middle. This “lost in the middle” suggests that these LLM-based models may not fully utilize the context provided by items that don’t occupy prominent positions in the input sequence (Liu et al., 2024).
There are a number of works on mitigating position bias, and we can categorize them into three lines based on our distribution alignment framework. (1) Prompting: This approach involves carefully designed prompts to encourage the models to disregard the input’s order (Hou et al., 2024; Dai et al., 2023). Nonetheless, due to LLMs’ prompt sensitivity, this requires precise and task-specific prompt engineering across various tasks and domains. (2) Data Augmentation: This approach has been explored in numerous studies, which is a form of data augmentation that involves random shuffling of candidates followed by aggregation to determine the final ranking. (Hou et al., 2024; Qin et al., 2023; Tang et al., 2023; Wu et al., 2023b; Luo et al., 2023; Zhang et al., 2023b). For instance, Tang et al. (2023) introduced a permutation self-consistency method, offering theoretical guarantees under certain conditions, enabling models to produce and aggregate potential outcomes from various candidate permutations, enhancing result stability. (3) Rebalancing: This method counteracts position bias by adjusting the prior distribution sensitive to positions. It recalibrates the model’s output, addressing the inherent bias towards item positions (Ma et al., 2023; Wu et al., 2023b).